Hache Corporation uses the weighted-average method in its process costing system. The first processing department, the Welding Department, started the month with 17,550 units in its beginning work in process inventory that were 20% complete with respect to conversion costs. The conversion cost in this beginning work in process inventory was $7,920. An additional 90,100 units were started into production during the month and 93,100 units were completed in the Welding Department and transferred to the next processing department. There were 14,550 units in the ending work in process inventory of the Welding Department that were 90% complete with respect to conversion costs. A total of $203,500 in conversion costs were incurred in the department during the month. The cost per equivalent unit for conversion costs for the month is closest to:

Multiple Choice

A. $2.020

B. $2.255

C. $1.991

D. $2.329
Answer #1

Ending units = 17550+90100-93100 = 14550

Equivalent units of conversion = 93100 + 14550*90% = 106195 units

Total cost = 203500 + 7920 = 211420

The cost per equivalent unit for conversion ccosts = $211420 / 106195 = $1.991

Option C. is correct answer.